OpenStudy (anonymous):

Which of the following represents the zeros of f(x) = x3 − 5x2 − 3x + 15? 5, square root of 3, −square root of 3 −5, −square root of 3, −square root of 3 5, −square root of 3, −square root of 3 −5, square root of 3, square root of 3

OpenStudy (jhannybean):

\[f(x) = x^3-5x^2-3x+15=0\]\[f(x) = x^2(x-5) +3(x-5) = 0\]\[f(x) = (x^2+3)(x-5)=0\]\[x^2+3=0 \qquad x=~?\]\[x-5=0 \qquad x=~?\]

OpenStudy (anonymous):

X=-3 X=5

OpenStudy (jhannybean):

You've got one right

OpenStudy (anonymous):

X^2=-3?

OpenStudy (jhannybean):

\[x^2+3=0 \implies x^2=-3\]\[x= \pm i\sqrt{3}\]Your answer choices are kind of weird.

OpenStudy (anonymous):

so its A then right

OpenStudy (jhannybean):

Oh I think Im the one that made the mistake.

OpenStudy (anonymous):

wait how

Parth (parthkohli):

yes it's x^2 - 3

OpenStudy (anonymous):

so it would still be A then right

Parth (parthkohli):

ya

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

do you mind helping me with one more?

OpenStudy (anonymous):

@Jhannybean @ParthKohli

OpenStudy (jhannybean):

correction: \((x^2-3)(x-5)=0\)

OpenStudy (jhannybean):

It wouldnt let me post. -_-

Parth (parthkohli):

i know that feel

OpenStudy (anonymous):

Which of the following represents the zeros of f(x) = x3 − 14x2 + 61x − 84? 7, 4, −3 7, 4, 3 7, −4, −3 7, −4, 3

OpenStudy (anonymous):

i believe it would be B

Parth (parthkohli):

Look at the choices. We know that \(7\) is definitely a zero. Divide the thing by \(x - 7\). The remaining is a quadratic polynomial that can be factored.

Parth (parthkohli):

or just plug in the choices lol idk who cares

Parth (parthkohli):

Vieta's Formulas @Jhannybean :)

OpenStudy (anonymous):

(x - 3) • (x - 4) • (x - 7)

OpenStudy (anonymous):

i got this

OpenStudy (anonymous):

so would it be B

OpenStudy (jhannybean):

=__=.... VF. How would that work here? we have one root = \((x-7)\) Then we'd find \(\alpha +\beta \) and \(\alpha\beta\) ?

OpenStudy (jhannybean):

@ParthKohli darn it you logged.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

srry i have to go

OpenStudy (jhannybean):

Np :)